ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 49,60,93 Yaroshevsky, Vassily, and Kuzmin, Victor, Central Aerohydrodynamic Institute, Russia Estimation of large deviations of dynamic system phase The approximate methods of probability estimation of such deviations occurence, including the problem of high level-crossing, are suggested and checked. These methods are based on the reduction of very small probability estimation nonlinear problem (P=1. e-6 -: - 1. e-8) to the solution of “equivalent” variational problem. Several model problems and particular problems of flight dynamics are considered (for example the problem of safe automatic landing).